Grade 11Physical Chemistry

Which of the following relates to light as a wave motion- diffraction, interference ,Photoelectric effect ,e is equal to MC square

Rajendra
8 Years agoGrade 11

1 Answer

Akashkumartiwari Akashkumartiwari
8 Years ago
Photoelectric effect.Grouping like terms:0 = (mD - (M + m)Δx) / MSolving for Δx:Δx = mD / (M + m)Since m is extremely small:Δx = mD / M (5)Substituting equation 5 into equation 4:(mDM) / M = ED / c2The mass of the box and D cancel out, leaving:m = E / c2Which rearranges